# Assessing Inf luential Observations in Pain Prediction using fMRI Data

### v2

@@ 4.2.2 Clustering-based MIP (ClusMIP) Detection Procedure
Motivated by [38, 39], one apparent solution to detect multiple influential observations on variable selection is to integrate the RGD scheme into the GDF metric (3), leading to the RGD-based GDF (RGDF) procedure provided in Section 8 of the Supplementary Materials Part I.
@@
To address the first limitation, an improved choice of the initial values for the KC algorithm is proposed by [1], leading to the so-called $K$-means++ algorithm (KC++). In addition, to resolve the challenges brought about by the high dimensionality, two categories of remedial proposals are available in the literature, which are known as the projection-based and regularized clustering. The projection-based scheme is a hybrid technique by first projecting the high-dimensional data onto a low-dimensional space, followed by partition via the KC. Two proposals under this category are the $t$-distributed stochastic neighbor embedding ($t$-SNE) [18] and spectral clustering [24], where the projected low-dimensional space for the later work is spanned by eigenvectors of the normalized graph Laplacians. In comparison, the regularized clustering augments a penalty term to the appropriate objective function to obtain estimates of penalized cluster centers. One such approach is called regularized KC [30], where a group LASSO penalty term is adopted and strong consistency of cluster centers is established [ibid., Theorem 1] for $p\rightarrow\infty$. These five clustering procedures, including the KC, KC++, $t$-SNE, spectral clustering and regularized KC are empirically compared on all the simulation configurations considered in Section 5. The corresponding graphical illustrations are reported in Section 10 of the Supplementary Materials Part I, where all of them exhibit analogously satisfactory detection of clusters. Indeed, in both finite and diverging dimensions, with a consistent clustering technique that is further refined via the CLT of the GDF measure (3) established in Theorem 2, the FPR of the foregoing ClusMIP algorithm is controlled at a given the FDR which is stated below in Theorem 3. The consistency conditions (B.1) to (B.2) [27], (C.1) to (C.7) [29], and the proof are given in Section 9 of the Supplementary Materials Part I:
@@ 5.1 Data Generation
Let $(n,p)=(100,1000)$ and $(160,480)$, where the latter dimension setting resembles those obsereved in the motivating fMRI study [21]. For each set of dimensions, $\textbf{X}\in\mathbb{R}^{n\times p}$ follows a zero-mean multivariate normal distribution, denoted $\textbf{X}\sim\mathcal{N}(\boldsymbol{0},\Sigma)$. Two possible variance-covariance matrices $\Sigma\in\mathbb{R}^{p\times p}$ are considered: the Toeplitz and the spatially correlated structures. Specifically, their $(i,j)$-entry is given by $\rho_{1}^{|i-j|}$ and $\rho_{2}^{|i-j|}\mathbbm{1}(i,j\in\{(p-S)/2+1,...,(p-S)/2+S\})$, $i=1,...,p$ and $j=1,...,p$, respectively. The spatially correlated formulation, where only the center $S$ predictors are dependent, mimics the brain connectivity patterns observed in fMRI. In our simulation study, we let $\rho_{1}\in\{0,0.5,0.8\}$, $\rho_{2}=0.8$ and $S=100$ so that four design matrices respectively denoted by $\textbf{X}_{\text{Tp(0)}}$, $\textbf{X}_{\text{Tp(0.5)}}$, $\textbf{X}_{\text{Tp(0.8)}}$ and $\textbf{X}_{\text{Sc(0.8)}}$ are generated for each dimension.
@@
Influential and clean observations are synthesized using the following two-step procedure: first, generate $n=100$ observations based on (1). Second, generate $n_{\text{infl}}$ points, denoted by $\{\textbf{Z}_{i,\text{infl}}=(\text{Y}_{i,\text{infl}},\textbf{X}_{i,\text{infl}}^{\top})^{\top}\in\mathbb{R}^{1001}:i=1,...,n_{\text{infl}}\}$, according to three perturbation schemes presented below, where $n_{\text{infl}}$ is the number of influential points. The first $n_{\text{infl}}$ points in the first step are then replaced by $\textbf{Z}_{i,\text{infl}},i=1,..,n_{\text{infl}}$, yielding the heterogeneous, contaminated dataset. Here, the three schemes respectively introduce contamination to the response vector only, design matrix only and both of them simultaneously. Moreover, the proportion of contamination $\zeta\coloneqq(n_{\text{infl}}/n)\times 100\%$ is set to be $5\%,10\%,15\%$ and $20\%$. The data generation is given in Section 11.1 of the Supplementary Materials Part I.
@@
    Scheme I (Contamination on the Response): $\textbf{X}_{i,\text{infl}}=\textbf{X}_{i}$ and $\text{Y}_{i,\text{infl}}=\text{Y}_{i_{\text{linear},\text{max}}}\\ +\epsilon_{i}+\kappa_{\text{linear},\textbf{Y}}$, where $\epsilon_{i}\sim N(0,1)$ and $\kappa_{\text{linear},\textbf{Y}}\in\{5,10,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}20}\}$.
@@
    Scheme II (Contamination on the Predictors): $\text{X}_{ij,\text{infl}}=\text{X}_{ij}+\kappa_{\text{linear},\textbf{X}}\cdot\mathbbm{1}\big{\{}j\in\{1,...,10\}\big{\}}$, $\kappa_{\text{linear},\textbf{X}}\in\{5,10,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}20}\}$, $\textbf{X}_{i,\text{infl}}=(\text{X}_{i1,\text{infl}},...,\text{X}_{ip,\text{infl}})^{\top}$, and $\text{Y}_{i,\text{infl}}=\text{Y}_{i}$.
@@
    Scheme III (Contamination on the Response and Predictors): $\text{X}_{ij,\text{infl}}=\\ \text{X}_{i_{\text{linear},\text{max}}j}+0.5\ \kappa_{\text{linear},\textbf{X}}\cdot\mathbbm{1}\big{\{}j\in\{1,...,10\}\big{\}}$, $\textbf{X}_{i,\text{infl}}=(\text{X}_{i1,\text{infl}},...,\text{X}_{ip,\text{infl}})^{\top}$, and $\text{Y}_{i,\text{infl}}=(\textbf{X}_{i,\text{infl}})^{\top}\boldsymbol{\beta}+\epsilon_{i}+\kappa_{\text{linear},\textbf{Y}}$, where $\epsilon_{i}\sim N(0,1)$ and $(\kappa_{\text{linear},\textbf{X}},\kappa_{\text{linear},\textbf{Y}})\in\{(5,5),(10,10),{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(20,20)}\}$.
@@ 5.2 Assessment Metrics
Two groups consisting of five metrics are examined. The first group contains the power ($\text{Power}_{\text{Detection}(\text{Selector})}$), false positive rate of detection ($\text{FPR}_{\text{Detection}(\text{Selector})}$), and computation time ($\text{Time}_{\text{Detection}(\text{Selector})}$). Here “Detection(Selector)” specifies the identification procedure “Detection” when a submodel is selected by “Selector”. They directly evaluate the detection performance. The second group includes the MSE of prediction and the empirical probability that the pre-specified true model is contained in the selected submodel (abbreviated as above), each computed both before and after removal of detected influential points for comparison purposes. They highlight the benefits of outlier detection on downstream inference. Here, since MIP has no model selection, the latter two metrics are excluded. These five metrics are formally defined in Section 11.2 of the Supplementary Materials Part I.
@@ 5.3 Methods and Implementation
Three detection methods are included: the ClusMIP, DF(LASSO) and MIP. In case of $\textbf{X}_{\text{Tp(0.5)}}$, $\textbf{X}_{\text{Tp(0.8)}}$ and $\textbf{X}_{\text{Sc(0.8)}}$, the ClusMIP is combined with LASSO, scaled LASSO (SLASSO), elastic net (ENET), SCAD and MCP, while the ENET is excluded for $\textbf{X}_{\text{Tp(0)}}$. Moreover, tuning parameter for all the regularized model selectors is chosen via the 10-fold cross-validation (CV), unless otherwise specified. A summary of methods and implementation is given in Section 11.3 of the Supplementary Materials Part I.
@@ 5.4 Simulation Results
Simulation results with distinctive patterns are presented. Others providing similar conclusion are in Sections 1 of the Supplementary Materials Part II for $(n,p)=(100,1000)$ and III for $(n,p)=(160,480)$, respectively.
@@
Figure 1: Perturbation Scheme I with $(n,p)=(100,1000)$,$\kappa_{\text{Linear},\textbf{Y}}=$5 and $\textbf{X}=\textbf{X}_{\text{Sc(0.8)}}$. Panel (A): power of detection; Panel (B): FPR of detection; Panels (C) and (D): MSE in prediction and probability that a selected submodel contains the true model before and after the removal of detected influential observations; Panel (E): algorithm running time in seconds.
@@
    Power: $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$ and $\text{Power}_{\text{MIP}}$ are generally higher (above 0.90) than the other procedures for the same contamination proportion $\zeta$ at moderate and high contamination scales $\kappa_{\text{linear},\textbf{Y}}$. This implies that SLASSO is sensitive to influential observations on variable selection, serving as a feasible candidate for use together with an outlier detection procedure. At low values of $\kappa_{\text{linear},\textbf{Y}}$, concerning under-performance was observed under certain scenarios for the MIP. For example, $\text{Power}_{\text{MIP}}$ drops from 0.20 to 0.10 as $\zeta$ increases from $15\%$ to $20\%$; see Panel (A) of Figure 1. In comparison, $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$ at $\zeta=15\%$ and $20\%$ consistently exceeds 0.97. Moreover, within the ClusMIP procedure, the detection powers of ClusMIP applied to LASSO, ENET, SCAD and MCP show higher variation and are generally lower than $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$, ranging between 0.45 and 0.89 under the particular simulation setting shown in Figure 1. In contrast, $\text{Power}_{\text{DF}(\text{LASSO})}$ (¡ 0.38 in Figure 1) is consistently among the lowest due to its leave-one-out design, indicating its limited ability to detect multiple influential points. In summary, ClusMIP(SLASSO) consistently exhibits better detection result than competing procedures, while under-performance of MIP is observed at low contamination intensity. Further, DF(LASSO) appers to be always outperformed in terms of detection.
@@
    MSE of Prediction: Lower prediction MSE after removing detected influential observations is consistently observed across all procedures, contamination proportions, and scales. For example, the MSE of ClusMIP(SLASSO) and DF(LASSO) declines from 35.94 to 2.84 and from 33.23 to 24.18, respectively, at $\zeta=20\%$ as seen in Panel (C) of Figure 1. While the extent of reduction in MSE varies according to the detection procedure, model selector, and contamination scales, MSE is always reduced after eliminating detected outliers, indicating the benefit of their detection. Moreover, low MSE after outlier removal is linked to improved detection performance. Noticeably, the MSE of SLASSO after the removal of outliers is always among the lowest, due to its strong detection performance. In comparison, the MSE after outlier removal for the remaining ClusMIP procedures showshigher fluctuation across simulation settings but generally corresponds to their detection power. For example, in Panels (A) and (C) of Figure 1, the MSE of ClusMIP(SCAD) after outlier removal (between 2.22 and 3.24) is the second lowest which agrees with the second-best position of its detection power performance (between 0.77 and 0.90). In contrast, the MSE of the DF(LASSO) after outlier removal is higher than that of the ClusMIP procedures due to its worse outlier detection.
@@
    Variable Selection: Improvement in the empirical probability that the true model is included in the selected submodel after removing the detected influential points is generally noted for all the procedures, indicating that these observations are indeed influential on variable selection. Similar to the reduction in MSE discussed above, the performance of this assessment metric hinges on the detection outcome. Due to reduction in the heterogeneity in the data, better detection performance contributes to higher probability of choosing a submodel containing the true model after outlier removal. This is analogous to results seen for detection power. This is numerically corroborated by comparing Panels (A) and (D) of Figure 1: ClusMIP(SLASSO) attains the highest detection power (0.97 – 0.99), followed by ClusMIP(SCAD) (0.77 – 0.90). On the other hand, while the selection probability is approximately 0 for both before outlier removal (i.e., based on the contaminated dataset), the two procedures achieve the highest selection probabilities after outlier removal, ranging between 0.90 – 1.00 and 0.66 – 0.86, respectively. In contrast, the selection probability after outlier removal for DF(LASSO) is 0 at $\zeta=10\%$, $15\%$ and $20\%$, which can be explained by its corresponding poor performance in detection.
@@
Figure 2: Perturbation Scheme II with (n,p)=(100,1000), $\kappa_{\text{Linear},\textbf{X}}$=10 and $\textbf{X}=\textbf{X}_{\text{Sc(0.8)}}$. Panel (A): power of detection; Panel (B): FPR of detection; Panels (C) and (D): MSE in prediction and probability that a selected submodel contains the true model before and after the removal of detected influential observations; Panel (E): algorithm running time in seconds.
@@
    FPR and Computation Time: they are similar to those under Perturbation Scheme I.

### v3

@@ 4.2.2 Clustering-based MIP (ClusMIP) Detection Procedure
Motivated by [38, 39], one apparent solution to detect multiple influential observations on variable selection is to integrate the RGD scheme into the GDF metric (3), leading to the RGD-based GDF (RGDF) procedure provided in Section 8 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)).
@@
To address the first limitation, an improved choice of the initial values for the KC algorithm is proposed by [1], leading to the so-called $K$-means++ algorithm (KC++). In addition, to resolve the challenges brought about by the high dimensionality, two categories of remedial proposals are available in the literature, which are known as the projection-based and regularized clustering. The projection-based scheme is a hybrid technique by first projecting the high-dimensional data onto a low-dimensional space, followed by partition via the KC. Two proposals under this category are the $t$-distributed stochastic neighbor embedding ($t$-SNE) [18] and spectral clustering [24], where the projected low-dimensional space for the later work is spanned by eigenvectors of the normalized graph Laplacians. In comparison, the regularized clustering augments a penalty term to the appropriate objective function to obtain estimates of penalized cluster centers. One such approach is called regularized KC [30], where a group LASSO penalty term is adopted and strong consistency of cluster centers is established [ibid., Theorem 1] for $p\rightarrow\infty$. These five clustering procedures, including the KC, KC++, $t$-SNE, spectral clustering and regularized KC are empirically compared on all the simulation configurations considered in Section 5. The corresponding graphical illustrations are reported in Section 10 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)), where all of them exhibit analogously satisfactory detection of clusters. Indeed, in both finite and diverging dimensions, with a consistent clustering technique that is further refined via the CLT of the GDF measure (3) established in Theorem 2, the FPR of the foregoing ClusMIP algorithm is controlled at a given the FDR which is stated below in Theorem 3. The consistency conditions (B.1) to (B.2) [27], (C.1) to (C.7) [29], and the proof are given in Section 9 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)):
@@ 5.1 Data Generation
Let $(n,p)=(100,1000)$ and $(160,480)$, where the latter dimension setting resembles those observed in the motivating fMRI study [21]. For each set of dimensions, $\textbf{X}\in\mathbb{R}^{n\times p}$ follows a zero-mean multivariate normal distribution, denoted $\textbf{X}\sim\mathcal{N}(\boldsymbol{0},\Sigma)$. Two possible variance-covariance matrices $\Sigma\in\mathbb{R}^{p\times p}$ are considered: the Toeplitz and the spatially correlated structures. Specifically, their $(i,j)$-entry is given by $\rho_{1}^{|i-j|}$ and $\rho_{2}^{|i-j|}\mathbbm{1}(i,j\in\{(p-S)/2+1,...,(p-S)/2+S\})$, i=1,,pi=1,...,p and j=1,,pj=1,...,p, respectively. The spatially correlated formulation, where only the center $S$ predictors are dependent, mimics the brain connectivity patterns observed in fMRI. In our simulation study, we let $\rho_{1}\in\{0,0.5,0.8\}$, $\rho_{2}=0.8$ and $S=100$ so that four design matrices respectively denoted by $\textbf{X}_{\text{Tp(0)}}$, $\textbf{X}_{\text{Tp(0.5)}}$, $\textbf{X}_{\text{Tp(0.8)}}$ and $\textbf{X}_{\text{Sc(0.8)}}$ are generated for each dimension.
@@
Influential and clean observations are synthesized using the following two-step procedure: first, generate $n=100$ observations based on (1). Second, generate $n_{\text{infl}}$ points, denoted by $\{\textbf{Z}_{i,\text{infl}}=(\text{Y}_{i,\text{infl}},\textbf{X}_{i,\text{infl}}^{\top})^{\top}\in\mathbb{R}^{1001}:i=1,...,n_{\text{infl}}\}$, according to three perturbation schemes presented below, where $n_{\text{infl}}$ is the number of influential points. The first $n_{\text{infl}}$ points in the first step are then replaced by $\textbf{Z}_{i,\text{infl}},i=1,..,n_{\text{infl}}$, yielding the heterogeneous, contaminated dataset. Here, the three schemes respectively introduce contamination to the response vector only, design matrix only and both of them simultaneously. Moreover, the proportion of contamination $\zeta\coloneqq(n_{\text{infl}}/n)\times 100\%$ is set to be $5\%,10\%,15\%$ and $20\%$. The data generation is given in Section 11.1 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)).
@@
    Scheme I (Contamination on the Response): $\textbf{X}_{i,\text{infl}}=\textbf{X}_{i}$ and $\text{Y}_{i,\text{infl}}=\text{Y}_{i_{\text{linear},\text{max}}}\\ +\epsilon_{i}+\kappa_{\text{linear},\textbf{Y}}$, where $\epsilon_{i}\sim N(0,1)$ and $\kappa_{\text{linear},\textbf{Y}}\in\{5,10,20\}$.
@@
    Scheme II (Contamination on the Predictors): $\text{X}_{ij,\text{infl}}=\text{X}_{ij}+\kappa_{\text{linear},\textbf{X}}\cdot\mathbbm{1}\big\{j\in\{1,...,10\}\big\}$, $\kappa_{\text{linear},\textbf{X}}\in\{5,10,20\}$, $\textbf{X}_{i,\text{infl}}=(\text{X}_{i1,\text{infl}},...,\text{X}_{ip,\text{infl}})^{\top}$, and $\text{Y}_{i,\text{infl}}=\text{Y}_{i}$.
@@
    Scheme III (Contamination on the Response and Predictors): $\text{X}_{ij,\text{infl}}=\\ \text{X}_{i_{\text{linear},\text{max}}j}+0.5\ \kappa_{\text{linear},\textbf{X}}\cdot\mathbbm{1}\big\{j\in\{1,...,10\}\big\}$, $\textbf{X}_{i,\text{infl}}=(\text{X}_{i1,\text{infl}},...,\text{X}_{ip,\text{infl}})^{\top}$, and $\text{Y}_{i,\text{infl}}=(\textbf{X}_{i,\text{infl}})^{\top}\boldsymbol{\beta}+\epsilon_{i}+\kappa_{\text{linear},\textbf{Y}}$, where $\epsilon_{i}\sim N(0,1)$ and $(\kappa_{\text{linear},\textbf{X}},\kappa_{\text{linear},\textbf{Y}})\in\{(5,5),(10,10),(20,20)\}$.
@@ 5.2 Assessment Metrics
Two groups consisting of six metrics were computed The first group consists of the power ($\text{Power}_{\text{Detection}(\text{Selector})}$), false positive rate of detection ($\text{FPR}_{\text{Detection}(\text{Selector})}$), mean number of detected influential points impacting the selection of stable/true signals/predictors ($\text{Stable}_{\text{Detection}(\text{Selector})}$), and computation time ($\text{Time}_{\text{Detection}(\text{Selector})}$). Specifically, $\text{Stable}_{\text{Detection}(\text{Selector})}$ determines whether each individual identified outlier alters the choice of at least two (out of a total three) pre-specified data-generating predictors. Here “Detection(Selector)” specifies the identification procedure “Detection” when a submodel is selected by “Selector”. They directly evaluate the detection performance. The second group includes the MSE of prediction and the empirical probability that the pre-specified true model is contained in the selected submodel (abbreviated as above), each computed both before and after removal of detected influential points for comparison purposes. They highlight the benefits of outlier detection on downstream inference. Here, since MIP has no model selection, the latter two metrics are excluded. These six metrics are formally defined in Section 11.2 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)).
@@ 5.3 Methods and Implementation
Three detection methods are included: the ClusMIP, DF(LASSO) and MIP. In case of $\textbf{X}_{\text{Tp(0.5)}}$, $\textbf{X}_{\text{Tp(0.8)}}$ and $\textbf{X}_{\text{Sc(0.8)}}$, the ClusMIP is combined with LASSO, scaled LASSO (SLASSO), elastic net (ENET), SCAD and MCP, while the ENET is excluded for $\textbf{X}_{\text{Tp(0)}}$. Moreover, tuning parameter for all the regularized model selectors is chosen via the 10-fold cross-validation (CV), unless otherwise specified. A summary of methods and implementation is given in Section 11.3 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)).
@@ 5.4 Simulation Results
Simulation results with distinctive patterns are presented. Others providing similar conclusion are in Sections 1 of the Supplementary Materials Part II (Zhang, Asgharian and Lindquist (2026)) for $(n,p)=(100,1000)$ and Part III for $(n,p)=(160,480)$, respectively.
@@
[…]
@@
    Power: $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$ and $\text{Power}_{\text{MIP}}$ are generally higher (above 0.90) than the other procedures for the same contamination proportion $\zeta$ at moderate and high contamination scales $\kappa_{\text{linear},\textbf{Y}}$. This implies that SLASSO is sensitive to influential observations on variable selection, serving as a feasible candidate for use together with an outlier detection procedure. At low values of $\kappa_{\text{linear},\textbf{Y}}$, concerning under-performance was observed under certain scenarios for the MIP. For example, $\text{Power}_{\text{MIP}}$ drops from 0.20 to 0.10 as $\zeta$ increases from $15\%$ to $20\%$; see Panel (A) of Figure 1. In comparison, $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$ at $\zeta=15\%$ and $20\%$ consistently exceeds 0.97. Moreover, within the ClusMIP procedure, the detection powers of ClusMIP applied to LASSO, ENET, SCAD and MCP show higher variation and are generally lower than $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$, ranging between 0.45 and 0.89 under the particular simulation setting shown in Figure 1. In contrast, $\text{Power}_{\text{DF}(\text{LASSO})}$ (¡ 0.38 in Figure 1) is consistently among the lowest due to its leave-one-out design, indicating its limited ability to detect multiple influential points. In summary, ClusMIP(SLASSO) consistently exhibits better detection result than competing procedures, while under-performance of MIP is observed at low contamination intensity. Further, DF(LASSO) appears to be always outperformed in terms of detection.
@@
    Stable Signal Selection: At each contamination rate $\zeta$, the ClusMIP of all model selectors detects significantly more outliers that impact the selection of true predictors than DF(LASSO)(Panel (C) of Figure 1). In other words, the ClusMIP detects a considerably higher number of genuine outliers, while few are identified by DF(LASSO). Specifically, $\text{Stable}_{\text{DF(LASSO)}}=1.95,1.4,2.75$ and $1.1$ at $\zeta=5\%,10\%,15\%$ and $20\%$ (i.e. at $n_{\text{infl}}=5,10,15$ and 20). In comparison, $\text{Stable}_{\text{ClusMIP(SLASSO)}}$ is $6.1,9.7,14.9$ and $18.95$. This result, corroborated by the pattern of detection power (Panel (A)), reinforces that ClusMIP outperforms DF(LASSO) in accurately identifying true outliers.
@@
    MSE of Prediction: Lower prediction MSE after removing detected influential observations is consistently observed across all procedures, contamination proportions, and scales. For example, the MSE of ClusMIP(SLASSO) and DF(LASSO) declines from 35.94 to 2.84 and from 33.23 to 24.18, respectively, at $\zeta=20\%$ as seen in Panel (E) of Figure 1. While the extent of reduction in MSE varies according to the detection procedure, model selector, and contamination scales, MSE is always reduced after eliminating detected outliers, indicating the benefit of their detection. Moreover, low MSE after outlier removal is linked to improved detection performance. Noticeably, the MSE of SLASSO after the removal of outliers is always among the lowest, due to its strong detection performance. In comparison, the MSE after outlier removal for the remaining ClusMIP procedures showshigher fluctuation across simulation settings but generally corresponds to their detection power. For example, in Panels (A) and (E) of Figure 1, the MSE of ClusMIP(SCAD) after outlier removal (between 2.22 and 3.24) is the second lowest which agrees with the second-best position of its detection power performance (between 0.77 and 0.90). In contrast, the MSE of the DF(LASSO) after outlier removal is higher than that of the ClusMIP procedures due to its worse outlier detection.
  6. 6.
    Variable Selection: Improvement in the empirical probability that the true model is included in the selected submodel after removing the detected influential points is generally noted for all the procedures, indicating that these observations are indeed influential on variable selection. Similar to the reduction in MSE discussed above, the performance of this assessment metric hinges on the detection outcome. Due to reduction in the heterogeneity in the data, better detection performance contributes to higher probability of choosing a submodel containing the true model after outlier removal. This is analogous to results seen for detection power. This is numerically corroborated by comparing Panels (A) and (F) of Figure 1: ClusMIP(SLASSO) attains the highest detection power (0.97 – 0.99), followed by ClusMIP(SCAD) (0.77 – 0.90). On the other hand, while the selection probability is approximately 0 for both before outlier removal (i.e., based on the contaminated dataset), the two procedures achieve the highest selection probabilities after outlier removal, ranging between 0.90 – 1.00 and 0.66 – 0.86, respectively. In contrast, the selection probability after outlier removal for DF(LASSO) is 0 at $\zeta=10\%$, $15\%$ and $20\%$, which can be explained by its corresponding poor performance in detection.
@@
Figure 2: Perturbation Scheme II with (n,p)=(100,1000), $\kappa_{\text{Linear},\textbf{X}}$=10 and $\textbf{X}=\textbf{X}_{\text{Sc(0.8)}}$: (A) and (B) power and FPR of detection; (C) average number of detected outliers that impact the selection of stable signals; (D) algorithm running time in seconds; (E) and (F) MSE in prediction and probability that a selected submodel contains the true model before and after removing detected influential observations.
@@
    FPR, Stable Signal Selection, and Computation Time: they are similar to those under Perturbation Scheme I.